OpenStudy (anonymous):

How does this model demonstrate the Pythagorean Theorem? A. The sum of the lengths of the shortest and the longest sides is equal to twice the length of the middle side. So double the length of the longer leg of any right triangle is equal to the sum of the shorter leg and the hypotenuse. B. The sum of the area of the two smaller squares is equal to the area of the larger square. So the sum of the lengths of the two legs of any right triangle squared is equal to the length of the hypotenuse squared.

OpenStudy (anonymous):

C. The sum of the area of the smallest and the largest squares is equal to the area of the middle square. So the sum of the lengths of the shorter leg and the hypotenuse of any right triangle squared is equal to the length of the middle leg squared. D. The length of the longest side minus two equals the length of the middle side. The length of the middle side minus two equals the length of the shortest side. So the length of the short leg of any right triangle is equal to the length of the middle leg minus 2, and the length of the hypotenuse is equal to the length of the middle leg plus 2.

OpenStudy (igreen):

It's obvious..

TheSmartOne (thesmartone):

@Preetha is the only QH on ._.

OpenStudy (igreen):

Pythagorean Theorem states that \(\sf a^2 + b^2 = c^2\) where 'a' and 'b' are the two legs and 'c' is the hypotenuse.

OpenStudy (igreen):

6 and 8 are the two legs, 10 is the hypotenuse.

OpenStudy (igreen):

Therefore: \(\sf 6^2 + 8^2 = 10^2\)

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

i have read the answers four times now, and have not found a suitable one

OpenStudy (misty1212):

this one The sum of the area of the two smaller squares is equal to the area of the larger square. So the sum of the lengths of the two legs of any right triangle squared is equal to the length of the hypotenuse squared. comes closest but it is not right

OpenStudy (misty1212):

it should be "the sum of the squares of the lengths of the two legs...."

OpenStudy (anonymous):

this one was just too odd. I wanted to see if anyone could help 0.o

TheSmartOne (thesmartone):